OpenStudy (studiousstudent):

Joan poured 7/8 liters of water into a bottle. It filled 5/6 of the bottle. How many liters of water will completely fill the bottle?

OpenStudy (anonymous):

5/6 x 6/5 = 1 So, 7/8 x 6/5 = 21/20

OpenStudy (anonymous):

Alright so to do this problem, you need to translate math into english. So 7/8 L of water is 5/6 of a bottle. so 7/8=5/6 (x) where x is the total volume of the bottle solve for x and get 21/20

OpenStudy (anonymous):

does this make sense?

OpenStudy (studiousstudent):

what do you mean by x is the total volume of the bottle... solve for x and get 21/20.

OpenStudy (studiousstudent):

Do you have to find the reciprocal?

OpenStudy (anonymous):

yes to solve the equation you multiply 6/5

OpenStudy (anonymous):

they tell you that 5/6 of the bottle is 7/8 litters

OpenStudy (studiousstudent):

So 7/8 x 6/5 = 21/20 = 1 1/20

OpenStudy (anonymous):

yes and g2g eat now if you need more help then become a fan

OpenStudy (studiousstudent):

1 1/20 as a mixed number

OpenStudy (studiousstudent):

Im a fan thanks

OpenStudy (studiousstudent):



OpenStudy (phi):

You could use ratios (but the fractions make it look ugly) \[ \frac{\frac{7}{8}}{\frac{5}{6}} = \frac{x}{1} \] that says: 7/8 liter is to 5/6 of a bottles as x (unknown liters) is to 1 bottle simplify: x/1 is x for the fraction: multiply top and bottom by 6/5 \[ x= \frac{\frac{7}{8}}{\frac{5}{6}}\cdot \frac{\frac{6}{5}}{\frac{6}{5}}\] the bottoms cancel and you are left with \[ x= \frac{7}{8}\cdot \frac{6}{5} = 42/40= 21/20= 1.05\]